OpenStudy (errinss):

A 12-foot ladder is leaning against a wall. The bottom of the ladder is 5 feet away from the bottom of the wall. Approximately how high up the wall does the top of the ladder reach? A. 2.4 feet B. 7.0 feet C. 13.0 feet D. 10.9 feet

OpenStudy (photon336):



rishavraj (rishavraj):

Apply Ppythogoras theorem :))

OpenStudy (errinss):



rishavraj (rishavraj):

yup :))

OpenStudy (errinss):

\[A^2+5^2=12^2\]

OpenStudy (errinss):

\[12^2-5^2=A^2\]

OpenStudy (errinss):

12^2-5^2 = 10.908

OpenStudy (errinss):

@Cheez

OpenStudy (dumbcow):

correct

rishavraj (rishavraj):

good work :)) @errinss

OpenStudy (errinss):

WOOHOO!!!!!